# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@ 5.2. The Kropina metric for chains on $S\Sigma$ and a variational characterization of conformal geodesics
$G^{\rm F}(\dot{\gamma}_{\epsilon},\dot{\gamma}_{\epsilon})$ $=G^{\rm F}\bigl{(}(R_{e^{\epsilon X}})_{*}\dot{\gamma},(R_{e^{\epsilon X}})_{*}\dot{\gamma}\bigr{)}$
$\quad+2\epsilon G^{\rm F}\bigl{(}(R_{e^{\epsilon X}})_{*}\dot{\gamma},\dot{X}^{\dagger}\bigr{)}+\epsilon^{2}G^{\rm F}\bigl{(}\dot{X}^{\dagger},\dot{X}^{\dagger}\bigr{)}$
$=G^{\rm F}(\dot{\gamma},\dot{\gamma})+2\epsilon G^{\rm F}\bigl{(}\dot{\gamma},({\rm Ad}(e^{\epsilon X})\dot{X})^{\dagger}\bigr{)}.$
@@
$\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}G^{\rm F}(\dot{\gamma}_{\epsilon},\dot{\gamma}_{\epsilon})=2G^{\rm F}(\dot{\gamma},\dot{X}^{\dagger})=2|\dot{x}|\omega^{3}(\dot{X}^{\dagger})=2|\dot{x}|\dot{X}^{3}.$
@@
$G^{\rm F}(K,\dot{\gamma}_{\epsilon})$ $=G^{\rm F}(Y^{\dagger},(R_{e^{\epsilon X}})_{*}\dot{\gamma}+\epsilon\dot{X}^{\dagger})$
$=G^{\rm F}\bigl{(}({\rm Ad}(e^{\epsilon X})Y)^{\dagger},\dot{\gamma}\bigr{)}$
$=\omega^{3}\bigl{(}({\rm Ad}(e^{\epsilon X})Y)^{\dagger}\bigr{)},$
@@
$\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}G^{\rm F}(K,\dot{\gamma}_{\epsilon})=\omega^{3}([X,Y]^{\dagger})=\omega^{3}\left(\begin{pmatrix}0&0&-X^{1}\\ 0&0&-X^{2}\\ X^{1}&X^{2}&0\end{pmatrix}^{\dagger}\right)=0.$
Thus, noting that $G^{F}(K,\dot{\gamma})=|\dot{x}|\omega^{3}(Y^{\dagger})=|\dot{x}|$, we have
[…]
Using this proposition, we prove the following variational characterization of conformal geodesics:
@@
Let $x(t)\ (a\leq t\leq b)$ be a regular curve on $(\Sigma,g)$ and $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$ its lift to $S\Sigma$. Let $F$ be the Kropina metric on $S\Sigma$ associated with a local section $s$ of $\mathcal{P}\to S\Sigma$ defined in a neighborhood of $\tilde{x}(t)$. Then, $x(t)$ is a conformal geodesic on $(\Sigma,g)$ if and only if it is a critical curve of the functional
[…]
@@
If $x(t)$ is a conformal geodesic, $\tilde{x}(t)$ is a chain on $S\Sigma$ by Theorem 5.4. Thus, $\tilde{x}(t)$ is a critical curve for the Kropina length and hence $x(t)$ is a critical curve of $\mathscr{F}$ for the variations fixing $\tilde{x}(a)$ and $\tilde{x}(b)$.
Conversely, suppose $x(t)$ is a critical curve of $\mathscr{F}$. We will show that the Kropina length functional $L$ is stationary under variations of $\tilde{x}(t)$. Let $y_{\epsilon}(t)$ be an arbitrary variation of $\tilde{x}(t)$ and let $V(t)=(\partial/\partial\epsilon)|_{\epsilon=0}y_{\epsilon}(t)$ be the variation vector field. We take a variation $x_{\epsilon}(t)$ of $x(t)$ which has the variation vector field $\pi_{*}V(t)$, where $\pi\colon S\Sigma\to\Sigma$ is the projection. Then, we can write as
@@
where $V_{1}(t):=(\partial/\partial\epsilon)|_{\epsilon=0}\tilde{x}_{\epsilon}(t)$ is the variation vector field of the lift $\tilde{x}_{\epsilon}(t):=(x_{\epsilon}(t),\dot{x}_{\epsilon}(t)/|\dot{x}_{\epsilon}(t)|)$ and $V_{2}(t)\in\mathrm{Ker}\,\pi_{*}$ is a vertical vector field. By the assumption, we have $L^{\prime}(V_{1})=\mathscr{F}^{\prime}(\pi_{*}V)=0$. Since we also have $L^{\prime}(V_{2})=0$ by Proposition 5.7, we obtain $L^{\prime}(V)=0$. ∎
5.3. The total torsion functional
We will discuss the relation between the functional $\mathscr{F}$ and the total torsion functional, which is defined as follows:
Let $x(t)\ (a\leq t\leq b)$ be a regular curve on $(\Sigma,g)$. We fix oriented orthonormal basis $A=(e_{1},e_{2},\dot{x}(a)/|\dot{x}(a)|)\in\mathcal{P}_{x(a)}$ and $B=(e^{\prime}_{1},e^{\prime}_{2},\dot{x}(b)/|\dot{x}(b)|)\in\mathcal{P}_{x(b)}$ at the endpoints. Then, we have the associated orientation preserving orthogonal map
$h\colon\dot{x}(a)^{\perp}\longrightarrow\dot{x}(b)^{\perp}$
@@ 5.3. The total torsion functional
for the parallel transport along $x(t)$ with respect to the normal connection $\nabla^{\perp}$. Then, the total torsion (functional) is defined by
$\mathscr{T}[x(t)]:=h^{-1}\circ P\in\mathrm{SO}(\dot{x}(a)^{\perp})\cong\mathbb{R}/2\pi\mathbb{Z}.$
@@
$g(\nabla_{\dot{x}}\nu,T\times\nu)=g(\nabla^{\perp}_{\dot{x}}\nu,T\times\nu)=-\dot{\alpha}.$
@@
We added the assumption $h(\nu(a))=\nu(b)$, which is missed in [7, Lemma 1]. We also modified the formula by putting the negative sign.
The formula (5.5) implies that the functional $\mathscr{T}$ is conformally invariant:
@@
Let $\mathscr{T}[x(t)]$ be the total torsion associated with orthonormal basis $A\in\mathcal{P}_{x(a)}$ and $B\in\mathcal{P}_{x(b)}$. Let $\widehat{g}=e^{2\Upsilon}g$ be a rescaled metric and let $\widehat{\mathscr{T}}[x(t)]$ be the total torsion for $\widehat{g}$ associated with the orthonormal basis $e^{-\Upsilon}A$ and $e^{-\Upsilon}B$. Then, we have
[…]
@@
The unit velocity vector field of $x(t)$ with respect to $\widehat{g}$ is given by $\widehat{T}:=e^{-\Upsilon}T$, and we can use $\widehat{\nu}:=e^{-\Upsilon}\nu$ as a unit normal vector field. Then, we have $\widehat{T}\,\widehat{\times}\,\widehat{\nu}=e^{-\Upsilon}T\times\nu$, where $\widehat{\times}$ denotes the cross product with respect to $\widehat{g}$. Moreover, the conformal transformation formula of the Levi-Civita connection gives
^$\widehat{\nabla}_{\dot{x}}\widehat{\nu}$ $=e^{-\Upsilon}\bigl{(}-(\partial_{t}\Upsilon)\nu+\nabla_{\dot{x}}\nu+(\partial_{t}\Upsilon)\nu+(\nu\cdot\Upsilon)\dot{x}\bigr{)}$
$=e^{-\Upsilon}\bigl{(}\nabla_{\dot{x}}\nu+(\nu\cdot\Upsilon)\dot{x}\bigr{)}.$
@@
$\widehat{g}(\widehat{\nabla}_{\dot{x}}\widehat{\nu},\widehat{T}\,\widehat{\times}\,\widehat{\nu})=g(\nabla_{\dot{x}}\nu,T\times\nu).$
The functionals $\mathscr{F}$ and $\mathscr{T}$ are related as follows:
@@
Let $\mathscr{F}$ be the functional given by (5.4). Let $\mathscr{T}$ be the total torsion functional associated to the orthonormal basis $s(\tilde{x}(a))$, $s(\tilde{x}(b))$ at the endpoints. Then, we have
[…]
@@
$-g(\nabla_{\dot{x}}\nu,T\times\nu)=-g(\nabla_{\dot{x}}\nu,\nu^{\prime})=\omega^{3}(s_{*}\dot{\tilde{x}}(t)).$
Hence we obtain the equation by (5.4), (5.5). ∎
It follows from this proposition and Theorem 5.8 that a conformal geodesic $x(t)\ (a\leq x\leq b)$ is characterized as a critical curve for the total torsion functional with fixed $A,B$ under the variations fixing $\tilde{x}(a)$, $\tilde{x}(b)$.
@@
[…]
@@
Chern–Knöppel–Pedit–Pinkall [7] calculated the variation of the total torsion functional:
We define $*R\in T^{*}\Sigma\otimes\mathfrak{so}(T\Sigma)$ by applying the Hodge star operator $*\colon\wedge^{2}T^{*}\Sigma\to T^{*}\Sigma$ to the curvature tensor $R\in\wedge^{2}T^{*}\Sigma\otimes\mathfrak{so}(T\Sigma)$. In the index notation, we have
@@
Theorem 5.13 ([7, Theorem 2]).
Let $x(t)\ (a\leq t\leq b)$ be a regular curve on $(\Sigma,g)$. Let $x_{\epsilon}(t)$ be a variation of $x(t)$ which fixes $(x(a),\dot{x}(a)/|\dot{x}(a)|)$ and $(x(b),\dot{x}(b)/|\dot{x}(b)|)$, and $V(t)=(\partial/\partial\epsilon)|_{\epsilon=0}x_{\epsilon}(t)$ its variation vector field. If we define the total torsion $\mathscr{T}[x_{\epsilon}(t)]$ by using common orthonormal basis at the endpoints, we have
$\frac{\partial}{\partial\epsilon}\Big{|}_{\epsilon=0}\mathscr{T}[x_{\epsilon}(t)]=\int_{a}^{b}g\bigl{(}T\times\nabla_{T}\nabla_{T}T+(*R)(T)T,V\bigr{)}|\dot{x}|dt,$
@@
We obtain an alternative proof of Theorem 1.2 from this formula: By (3.1), we have
$(*R)_{mkl}=\frac{1}{2}\varepsilon_{m}{}^{ij}R_{ijkl}=\frac{1}{2}(\varepsilon_{m}{}^{i}{}_{l}P_{ik}-\varepsilon_{ml}{}^{j}P_{jk}+\varepsilon_{mk}{}^{j}P_{jl}-\varepsilon_{m}{}^{i}{}_{k}P_{il})$
@@
$\bigl{(}(*R)(T)T\bigr{)}^{k}$ $=(*R)_{m}{}^{k}{}_{l}T^{m}T^{l}$
$=\frac{1}{2}(\varepsilon_{m}{}^{k}{}_{j}P^{j}{}_{l}T^{m}T^{l}-\varepsilon_{mi}{}^{k}P^{i}{}_{l}T^{m}T^{l})$
$=-\bigl{(}T\times P(T)\bigr{)}^{k}.$
@@
$T\times\nabla_{T}\nabla_{T}T+(*R)(T)T=T\times\bigl{(}\nabla_{T}\nabla_{T}T-P(T)\bigr{)},$

### v4

@@ 5.2 The Kropina metric for chains on $\boldsymbol{S\Sigma}$ and a variational
$G^{\rm F}(\dot{\gamma}_{\epsilon},\dot{\gamma}_{\epsilon})={}$ $G^{\rm F}((R_{{\rm e}^{\epsilon X}})_{*}\dot{\gamma},(R_{{\rm e}^{\epsilon X}})_{*}\dot{\gamma})2\epsilon G^{\rm F}\bigl((R_{{\rm e}^{\epsilon X}})_{*}\dot{\gamma},\dot{X}^{\dagger}\bigr)+\epsilon^{2}G^{\rm F}\bigl(\dot{X}^{\dagger},\dot{X}^{\dagger}\bigr)$
$={}$ $G^{\rm F}(\dot{\gamma},\dot{\gamma})+2\epsilon G^{\rm F}\bigl(\dot{\gamma},\bigl({\rm Ad}\bigl({\rm e}^{\epsilon X}\bigr)\dot{X}\bigr)^{\dagger}\bigr).$
@@
$\frac{\partial}{\partial\epsilon}\bigg|_{\epsilon=0}G^{\rm F}(\dot{\gamma}_{\epsilon},\dot{\gamma}_{\epsilon})=2G^{\rm F}\bigl(\dot{\gamma},\dot{X}^{\dagger}\bigr)=2|\dot{x}|\omega^{3}\bigl(\dot{X}^{\dagger}\bigr)=2|\dot{x}|\dot{X}^{3}.$
@@
$G^{\rm F}(K,\dot{\gamma}_{\epsilon})=G^{\rm F}\bigl(Y^{\dagger},(R_{{\rm e}^{\epsilon X}})_{*}\dot{\gamma}+\epsilon\dot{X}^{\dagger}\bigr)=G^{\rm F}\bigl(\bigl({\rm Ad}\bigl({\rm e}^{\epsilon X}\bigr)Y\bigr)^{\dagger},\dot{\gamma}\bigr)=\omega^{3}\bigl(\bigl({\rm Ad}\bigl({\rm e}^{\epsilon X}\bigr)Y\bigr)^{\dagger}\bigr),$
@@
$\frac{\partial}{\partial\epsilon}\bigg|_{\epsilon=0}G^{\rm F}(K,\dot{\gamma}_{\epsilon})=\omega^{3}\bigl([X,Y]^{\dagger}\bigr)=\omega^{3}\left(\begin{pmatrix}0&0&-X^{1}\\ 0&0&-X^{2}\\ X^{1}&X^{2}&\hphantom{-}0\end{pmatrix}^{\dagger}\right)=0.$
Thus, noting that $G^{F}(K,\dot{\gamma})=|\dot{x}|\omega^{3}\bigl(Y^{\dagger}\bigr)=|\dot{x}|$, we have
$\frac{\mathrm{d}}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}\int_{a}^{b}\frac{G^{\rm F}(\dot{\gamma}_{\epsilon}(t),\dot{\gamma}_{\epsilon}(t))}{G^{\rm F}(K,\dot{\gamma}_{\epsilon}(t))}\,{\rm d}t=2\int_{a}^{b}\dot{X}^{3}(t)\,{\rm d}t=2\bigl(X^{3}(b)-X^{3}(a)\bigr)=0.$
Using this proposition, we prove the following variational characterization of conformal geodesics.
@@
Let $x(t)$, $a\leq t\leq b$, be a regular curve on $(\Sigma,g)$ and $\tilde{x}(t)=(x(t),\dot{x}(t)/|\dot{x}(t)|)$ its lift to $S\Sigma$. Let $F$ be the Kropina metric on $S\Sigma$ associated with a local section $s$ of $\mathcal{P}\to S\Sigma$ defined in a neighborhood of $\tilde{x}(t)$. Then, $x(t)$ is a conformal geodesic on $(\Sigma,g)$ if and only if it is a critical curve of the functional
$\mathscr{L}[x(t)]=\int_{a}^{b}F\bigl(\tilde{x}(t),\dot{\tilde{x}}(t)\bigr)\,{\rm d}t\biggl(=2\int_{a}^{b}\omega^{3}\bigl(s_{*}\dot{\tilde{x}}\bigr)\,{\rm d}t\biggr)$ (5.4)
@@
If $x(t)$ is a conformal geodesic, $\tilde{x}(t)$ is a chain on $S\Sigma$ by Theorem 5.4. Thus, $\tilde{x}(t)$ is a critical curve for the Kropina length and hence $x(t)$ is a critical curve of $\mathscr{L}$ for the variations fixing $\tilde{x}(a)$ and $\tilde{x}(b)$.
Conversely, suppose $x(t)$ is a critical curve of $\mathscr{L}$. We will show that the Kropina length functional $L$ is stationary under variations of $\tilde{x}(t)$. Let $y_{\epsilon}(t)$ be an arbitrary variation of $\tilde{x}(t)$ and let $V(t)=(\partial/\partial\epsilon)|_{\epsilon=0}y_{\epsilon}(t)$ be the variation vector field. We take a variation $x_{\epsilon}(t)$ of $x(t)$ which has the variation vector field $\pi_{*}V(t)$, where $\pi\colon S\Sigma\to\Sigma$ is the projection. Then, we can write as
@@
where $V_{1}(t):=(\partial/\partial\epsilon)|_{\epsilon=0}\tilde{x}_{\epsilon}(t)$ is the variation vector field of the lift $\tilde{x}_{\epsilon}(t):=(x_{\epsilon}(t),\dot{x}_{\epsilon}(t)/|\dot{x}_{\epsilon}(t)|)$ and $V_{2}(t)\in\operatorname{Ker}\pi_{*}$ is a vertical vector field. By the assumption, we have $L^{\prime}(V_{1})=\mathscr{L}^{\prime}(\pi_{*}V)=0$. Since we also have $L^{\prime}(V_{2})=0$ by Proposition 5.7, we obtain $L^{\prime}(V)=0$. ∎
5.3 The total torsion functional
We will discuss the relation between the functional $\mathscr{L}$ and the total torsion functional, which is defined as follows.
Let $x(t)$, $a\leq t\leq b$, be a regular curve on $(\Sigma,g)$. We fix oriented orthonormal basis $A=(e_{1},e_{2},\dot{x}(a)/|\dot{x}(a)|)\in\mathcal{P}_{x(a)}$ and $B=\bigl(e^{\prime}_{1},e^{\prime}_{2},\dot{x}(b)/|\dot{x}(b)|\bigr)\in\mathcal{P}_{x(b)}$ at the endpoints. Then, we have the associated orientation preserving orthogonal map
$h\colon\ \dot{x}(a)^{\perp}\longrightarrow\dot{x}(b)^{\perp},$
@@ 5.3 The total torsion functional
for the parallel transport along $x(t)$ with respect to the normal connection $\nabla^{\perp}$. Then, the total torsion ((functional)) is defined by
$\mathscr{T}[x(t)]:=h^{-1}\circ P\in\mathrm{SO}\bigl(\dot{x}(a)^{\perp}\bigr)\cong\mathbb{R}/2\pi\mathbb{Z}.$
@@
$g(\nabla_{\dot{x}}\nu,T\times\nu)=g\bigl(\nabla^{\perp}_{\dot{x}}\nu,T\times\nu\bigr)=-\dot{\alpha}.$
@@
We added the assumption $h(\nu(a))=\nu(b)$, which is missed in [10, Lemma 1]. We also modified the formula by putting the negative sign.
The formula (5.5) implies that the functional $\mathscr{T}$ is conformally invariant.
@@
Let $\mathscr{T}[x(t)]$ be the total torsion associated with orthonormal basis $A\in\mathcal{P}_{x(a)}$ and $B\in\mathcal{P}_{x(b)}$. Let $\widehat{g}={\rm e}^{2\Upsilon}g$ be a rescaled metric and let $\widehat{\mathscr{T}}[x(t)]$ be the total torsion for $\widehat{g}$ associated with the orthonormal basis ${\rm e}^{-\Upsilon}A$ and ${\rm e}^{-\Upsilon}B$. Then, we have
$\mathscr{T}[x(t)]\equiv\widehat{\mathscr{T}}[x(t)]\mod 2\pi\mathbb{Z}.$
@@
The unit velocity vector field of $x(t)$ with respect to $\widehat{g}$ is given by $\widehat{T}:={\rm e}^{-\Upsilon}T$, and we can use $\widehat{\nu}:={\rm e}^{-\Upsilon}\nu$ as a unit normal vector field. Then, we have $\widehat{T}\,\widehat{\times}\,\widehat{\nu}={\rm e}^{-\Upsilon}T\times\nu$, where $\widehat{\times}$ denotes the cross product with respect to $\widehat{g}$. Moreover, the conformal transformation formula of the Levi-Civita connection gives
$\widehat{\nabla}_{\dot{x}}\widehat{\nu}={\rm e}^{-\Upsilon}(-(\partial_{t}\Upsilon)\nu+\nabla_{\dot{x}}\nu+(\partial_{t}\Upsilon)\nu+(\nu\cdot\Upsilon)\dot{x})={\rm e}^{-\Upsilon}(\nabla_{\dot{x}}\nu+(\nu\cdot\Upsilon)\dot{x}).$
@@
$\widehat{g}\bigl(\widehat{\nabla}_{\dot{x}}\widehat{\nu},\widehat{T}\,\widehat{\times}\,\widehat{\nu}\bigr)=g\bigl(\nabla_{\dot{x}}\nu,T\times\nu\bigr).$
The functionals $\mathscr{L}$ and $\mathscr{T}$ are related as follows.
@@
Let $\mathscr{L}$ be the functional given by (5.4). Let $\mathscr{T}$ be the total torsion functional associated to the orthonormal basis $s(\tilde{x}(a))$, $s(\tilde{x}(b))$ at the endpoints. Then, we have
$\mathscr{L}[x(t)]\equiv 2\mathscr{T}[x(t)]\mod 2\pi\mathbb{Z}.$
@@
$g(\nabla_{\dot{x}}\nu,T\times\nu)=-g\bigl(\nabla_{\dot{x}}\nu,\nu^{\prime}\bigr)=\omega^{3}\bigl(s_{*}\dot{\tilde{x}}(t)\bigr).$
Hence we obtain the equation by (5.4) and (5.5). ∎
It follows from this proposition and Theorem 5.8 that a conformal geodesic $x(t)$, $a\leq x\leq b$, is characterized as a critical curve for the total torsion functional with fixed $A$, $B$ under the variations fixing $\tilde{x}(a)$, $\tilde{x}(b)$.
@@
$\mathscr{T}[x(t)]\equiv-\int_{a}^{b}\frac{\det(\dot{x},\nabla_{\dot{x}}\dot{x},\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x})}{|\dot{x}\times\nabla_{\dot{x}}\dot{x}|^{2}}|\dot{x}|\,{\rm d}t+c\mod 2\pi\mathbb{Z}$
@@
Chern–Knöppel–Pedit–Pinkall [10] calculated the variation of the total torsion functional. We define $R\in T^{*}\Sigma\otimes\mathfrak{so}(T\Sigma)$ by applying the Hodge star operator $\colon\wedge^{2}T^{*}\Sigma\to T^{*}\Sigma$ to the curvature tensor $R\in\wedge^{2}T^{*}\Sigma\otimes\mathfrak{so}(T\Sigma)$. In the index notation, we have
@@
Theorem 5.13 ([10, Theorem 2]).
Let $x(t)$, $a\leq t\leq b$, be a regular curve on $(\Sigma,g)$. Assume $x_{\epsilon}(t)$ is a variation of $x(t)$ which fixes $(x(a),\dot{x}(a)/|\dot{x}(a)|)$ and $(x(b),\dot{x}(b)/|\dot{x}(b)|)$, and $V(t)=(\partial/\partial\epsilon)|_{\epsilon=0}x_{\epsilon}(t)$ its variation vector field. If we define the total torsion $\mathscr{T}[x_{\epsilon}(t)]$ by using common orthonormal basis at the endpoints, we have
$\frac{\mathrm{d}}{\mathrm{d}\epsilon}\bigg|_{\epsilon=0}\mathscr{T}[x_{\epsilon}(t)]=\int_{a}^{b}g\bigl(T\times\nabla_{T}\nabla_{T}T+(*R)(T)T,V\bigr)|\dot{x}|\,{\rm d}t,$
@@
We obtain an alternative proof of Theorem 1.2 from this formula. By (3.1), we have
$(*R)_{mkl}=\frac{1}{2}\varepsilon_{m}{}^{ij}R_{ijkl}=\frac{1}{2}\bigl(\varepsilon_{m}{}^{i}{}_{l}P_{ik}-\varepsilon_{ml}{}^{j}P_{jk}+\varepsilon_{mk}{}^{j}P_{jl}-\varepsilon_{m}{}^{i}{}_{k}P_{il}\bigr)$
@@
$((*R)(T)T)^{k}=(*R)_{m}{}^{k}{}_{l}T^{m}T^{l}=\frac{1}{2}\bigl(\varepsilon_{m}{}^{k}{}_{j}P^{j}{}_{l}T^{m}T^{l}-\varepsilon_{mi}{}^{k}P^{i}{}_{l}T^{m}T^{l}\bigr)=-(T\times P(T))^{k}.$
@@
$T\times\nabla_{T}\nabla_{T}T+(*R)(T)T=T\times(\nabla_{T}\nabla_{T}T-P(T)),$
@@
Acknowledgements
The author thanks the anonymous referees for their comments. This work was partially supported by JSPS KAKENHI Grant Number 22K13922.